# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

### v2

@@ Abstract
Benefiting from contrastively trained visual encoders on large-scale natural scene images, Large Multimodal Models (LMMs) have achieved remarkable performance across various visual perception tasks. However, the inherent limitations of contrastive learning upon summarized descriptions fundamentally restrict the capabilities of models in meticulous reasoning, particularly in crucial scenarios of geometric problem-solving. To enhance geometric understanding, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train CLIP using our hard negative learning method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations. These analyses yield important insights into optimizing hard negative strategies for geometric reasoning tasks. https://github.com/THU-KEG/MMGeoLM.
Introduction
@@
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic 2024; OpenAI 2023; Bai et al. 2023). It requires fine-grained identification of visual elements (Lu et al. 2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al. 2024).
@@
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al. 2024; Yang et al. 2023; Goel et al. 2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI 2024a), Claude-3 (Anthropic 2024), and Qwen2.5-VL (Bai et al. 2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
[…]
@@
To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al. 2023; Zhang et al. 2024b; Peng et al. 2024a, b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al. 2024; Wang et al. 2024) by aligning images with corresponding captions. However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al. 2024a; Doveh et al. 2023a). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
@@
AltCLIP In-Batch Randomly-sampled Negative (400K) 24.6
AltCLIP MMCLIP ①: Retrieval-based Negative (100K) 26.6
AltCLIP MMCLIP ②: Rule-based Negative (100K) 28.1
AltCLIP MMCLIP ③: Image-based Negative (4K) 28.2
@@
Method
Negative Images Construction
We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives. Existing generation methods (Zhang et al. 2024b; Zou et al. 2024; Wei et al. 2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic geometry problems (Sun et al. 2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock, Jiang, and July 2002). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al. 2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind 2025). Upon evaluating 123 generated captions, we find that the accuracy of captions in representing geometric elements reaches 100%.
To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints.
Negative Captions Construction
@@
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al. 2020; Xiong et al. 2020; Huang et al. 2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al. 2020; Zhou et al. 2022; Yang et al. 2024). To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al. 2024b), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS (Zhou et al. 2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
@@
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al. 2022; Zhang, Awal, and Agrawal 2024), while others leverage in-context learning to generate meaningful modifications (Patel et al. 2025) . However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al. 2024) and identify four major types of image element recognition errors:
@@
To address these error types, we design rule-based strategies that use GLM-4 (GLM et al. 2024) to modify each positive caption and generate 10 hard negative examples.
These text-based negatives are used in model training, and two representative types are shown in Figure 2.
Hard Negative CLIP Training
@@
Though CLIP-trained vision encoders are widely used in LMMs (Radford et al. 2021; Doveh et al. 2023b), they typically combine random examples for a batch, which limits their capacity for fine-grained image understanding during training. While methods like NegCLIP (Yuksekgonul et al. 2022) introduce hard negatives, they still rely heavily on randomly selected negatives within a batch. Considering the methods in the above section have allowed us to construct a large number of hard negatives, we propose a MMCLIP training strategy, which focuses on each single sample and only uses its corresponding hard negatives for training. In this way, the vision encoder is forced to distinguish subtle differences, thereby enhancing its fine-grained geometry understanding capacities.
Formally, given a text-based negative training dataset $D=\{(I_{i},T_{i}^{+},\{T_{j}^{-}\}_{j=1}^{N})\}_{i=1}^{M}$, where $I_{i}$ represents the image, $T_{i}^{+}$ denotes the positive samples and $T_{j}^{-}$ is the negative samples conducted around the image. Our loss function is formulated as follows:
$\mathcal{L}\!=\!\frac{1}{M}\sum_{i=1}^{M}\left[\!-\!\log\frac{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}}{\exp\big{(}s(I_{i},T_{i}^{+})\big{)}\!+\!\sum\limits_{j=1}^{N}\exp\big{(}s(I_{i},T_{j}^{-})}\big{)}\right]$ (1)
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings. Since negative samples are constructed around elements within the image, their score $s(I_{i},T_{j}^{-})$ is expected to be higher than a randomly chosen one. From the above equation, we calculate the gradient for positive and negative samples:
$\frac{\partial\mathcal{L}}{\partial s_{i}^{+}}=-\frac{\sum_{j=1}^{N}\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (2)
$\frac{\partial\mathcal{L}}{\partial s^{-}_{j}}=\frac{\exp(s^{-}_{j})}{\exp(s_{i}^{+})+\sum_{j=1}^{N}\exp(s^{-}_{j})}$ (3)
where $s_{i}^{+}=s(I_{i},T_{i}^{+})$ and $s_{j}^{-}=s(I_{i},T_{j}^{-})$. Since the number of hard negatives is unrestricted, the gradient function guided by Equation 3 can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2. Unlike in-batch training, which relies on batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for rule-based or retrieval-based unimodal negatives. We further validate different training strategies only for hard negatives in the ablation experiments.
Supervised Fine-tuning Data
@@
Multimodal Mathematical Training
Architecture
We adopt the LLaVA architecture (Liu et al. 2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use MAmmoTH2-7B (Yue et al. 2024) and Qwen2.5-7B-Instruct (Qwen et al. 2025). The vision encoder is based on AltCLIP (Chen et al. 2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
Training Details
@@
Our training pipeline, illustrated in Figure 3, consists of three stages covering both main and ablation experiments. In the first stage, we perform vision-text alignment training. For text-based hard negatives, we use the Mavis image-caption alignment dataset (Zhang et al. 2024b). For image-based negatives, we collect 4K geometry questions from real middle-school exams. Additional details, including the types of negatives and the negative-to-positive ratio, are provided in Figure 3. In the second stage, we tune the MLP adapter using 67K image-text pairs from the G-LLaVA dataset (Gao et al. 2023). In the third stage, we conduct supervised fine-tuning. For main experiments, we use a combined dataset comprising 300K Mavis instruction data (Zhang et al. 2024b), 117K G-LLaVA instruction data (Gao et al. 2023), and 17K open-ended geometry questions. For ablation studies, we only use the 17K newly collected geometry problems for efficient validation. We name the final trained model as MMGeoLM.
Experiments
LMM Geometric Reasoning Evaluation
Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA (Chen et al. 2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al. 2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al. 2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-Math (Sun et al. 2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic 2024), Claude-3.5-Sonnet, GPT-4o (OpenAI 2024a), and GPT-4V. Open-source LMMs: MAVIS-7B (Zhang et al. 2024b), G-LLaVA-7B (Gao et al. 2023), InternVL2-8B (Chen et al. 2024), LLaVA-OneVision-7B (Li et al. 2024), Qwen2.5-VL-7B-Instruct (Bai et al. 2025), Chimera-Reasoner-8B (Peng et al. 2024b), InternLM-XComposer2-7B (Dong et al. 2024), Phi-3-Vision-128K-Instruct (Abdin et al. 2024), and Math-LLaVA-13B (Shi et al. 2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al. 2021; Lu et al. 2023). Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al. 2024b; Sun et al. 2024).
Overall Results
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA (Peng et al. 2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-MAmmoTH2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-MAmmoTH2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-Math benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the w/o MMCLIP (AltCLIP trained only with random negatives in a batch), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
Ablation Study 1: Different Negatives Types

### v3

# MMGeoLM: Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

@@ Abstract
Large Multimodal Models (LMMs) typically build on ViTs (e.g., CLIP), yet their training with simple random in-batch negatives limits the ability to capture fine-grained visual differences, particularly in geometric scenarios. To address this challenge, we propose a novel hard negative contrastive learning framework for the vision encoder, which combines image-based contrastive learning using generation-based hard negatives created by perturbing diagram generation code, and text-based contrastive learning using rule-based negatives derived from modified geometric descriptions and retrieval-based negatives selected based on caption similarity. We train a vision encoder (CLIP) using our hard negative training method, namely MMCLIP (Multimodal Math CLIP), and subsequently train an LMM for geometric problem-solving. Experiments show that our trained model, MMGeoLM, significantly outperforms other open-source models on three geometric reasoning benchmarks. Even with a size of 7B, it can rival powerful closed-source models like GPT-4o. We further conduct ablation studies to analyze three key factors: hard negative types, the efficiency of image-based negatives, and training configurations. These analyses yield important insights into optimizing the training pipeline of vision encoder for fine-grained geometric reasoning tasks. https://github.com/THU-KEG/MMGeoLM.
11footnotetext: Equal Contribution.
1 Introduction
@@
Geometric mathematical reasoning has garnered significant attention as an essential capability for large multimodal models (Anthropic, 2024; OpenAI, 2023; Bai et al., 2023). It requires fine-grained identification of visual elements (Lu et al., 2023) within the given images, such as geometric shapes, spatial configurations, and the relationships between them (He et al., 2024).
@@
However, the “eyes” of most existing LMMs, i.e., their pretrained vision encoders such as CLIP (Patel et al., 2024; Yang et al., 2023; Goel et al., 2022), are primarily trained on general visual datasets that do not emphasize the intricate features necessary for specialized mathematical reasoning. Therefore, these models often fail to understand the nuanced geometric information accurately and produce incorrect reasoning and answers. As shown in Figure 1, facing a simple parallel line problem, the leading LMMs such as GPT-4o (OpenAI, 2024a), Claude-3 (Anthropic, 2024), and Qwen2.5-VL (Bai et al., 2025) all hallucinate non-existent elements or misinterpret spatial relationships (e.g., $\triangle ABC$, $\triangle ABE$, and the concept of similar triangles), exhibiting notable deficiencies in capturing the intricate geometric details.
@@
To address these shortcomings, recent research has focused on strategies such as fine-tuning on specialized mathematical datasets (Gao et al., 2023; Zhang et al., 2024c; Peng et al., 2024a; b) or utilizing massive image-caption pairs to enhance the visual perception ability of LMMs (Qi et al., 2024; Wang et al., 2024) by aligning images with corresponding captions. However, relying solely on positive image-caption pairs may lead to spurious alignment, with models may behave like “bag of words” rather than achieving semantic understanding (Zhang et al., 2024b; Doveh et al., 2023a). To achieve more robust and semantically meaningful alignment, it is crucial to incorporate hard negative samples—semantically similar but mismatched pairs—which push the vision encoder to learn finer distinctions beyond shallow correlations. Therefore, to further improve LMMs’ abilities for capturing geometric information, we investigate a key question: How can we systematically construct hard negatives tailored for geometric reasoning?
Figure 1: Examples of hallucination: non-existent elements and relation misinterpretation.
@@
AltCLIP In-Batch Randomly-sampled Neg (400K) 24.6
AltCLIP MMCLIP ①: Retrieval Neg (100K) 26.6
AltCLIP MMCLIP ②: Rule Neg (100K) 28.1
AltCLIP MMCLIP ③: Image Neg (4K) 28.2
@@
2 Hard Negative Data Construction
2.1 Negative Images Construction
We construct negative geometric examples by generating geometric diagram code for positive images and perturbing it to create negatives. Existing generation methods (Zhang et al., 2024c; Zou et al., 2024; Wei et al., 2024) rely on handcrafted rules that cover limited elements and overlook inter-element relationships, resulting in a gap from real-world problems. In contrast, authentic geometry problems (Sun et al., 2024) are more diverse and accurate, with geometric elements described either explicitly (e.g., “AB is 8 cm”) or implicitly through reasoning (e.g., using the Pythagorean theorem), and reflected in the corresponding geometric diagrams (McClintock et al., 2002). Benefiting from advancements in mathematical reasoning and code-generation capabilities of LLMs (OpenAI, 2024b), which can effectively parse geometric conditions, perform necessary reasoning, and generate executable code to produce geometric diagrams closely matching the original images. As illustrated in Figure 2, providing the LLM with both the problem and the answer yields code and a caption. Executing the code produces diagrams that not only replicate the original figures but may also include additional numeric markings.
To handle occasional errors in model-generated code, we adopt a lightweight model (GLM et al., 2024) to correct syntax issues when the code fails. Subsequently, we generate detailed captions directly from code using Gemini 2.5 Pro (DeepMind, 2025). Upon evaluating 123 generated captions, three independent annotators confirmed that the accuracy of captions in representing geometric elements reaches 100%. Further details are provided in Appendix F.
To create the negative images, we design prompts for LLMs to generate 10 negative captions for each positive caption, differing subtly but critically. We then employ Gemini 2.5 Pro to modify the Python scripts based on these perturbed captions. These modified scripts produce diagrams that closely resemble the originals while aligning with the intended constraints. Representative examples of the constructed image-based negatives are provided in Appendix B.
2.2 Negative Captions Construction
@@
Many previous works in open-domain QA take the top-ranked instances recalled by the retriever as negative examples to further improve retriever performance (Karpukhin et al., 2020; Xiong et al., 2020; Huang et al., 2020). However, they suffer from the risks of introducing false negatives, which means some related instances are incorrectly treated as negatives (Xiong et al., 2020; Zhou et al., 2022; Yang et al., 2024). To address this issue, we build upon the recently proposed Mavis dataset (Zhang et al., 2024c), which mitigates false negatives by ensuring a one-to-one correspondence between each image and its caption. We use SimANS (Zhou et al., 2022) to encode all captions, compute pairwise similarities, and retrieve the top 100 most similar captions per image as hard negatives.
@@
Existing studies modify positive captions by randomly altering nouns, attributes, and relationships (Yuksekgonul et al., 2022; Zhang et al., 2024a), while others leverage in-context learning to generate meaningful modifications (Patel et al., 2025). However, these methods either introduce semantic ambiguity or focus solely on the vision encoder level, without considering downstream reasoning errors, as exemplified by the misunderstanding in Figure 1, where parallel lines are mistaken as forming a triangle.
To address this issue, we analyze the evaluation results of LMMs on MM-Math (Sun et al., 2024) and identify four major types of image element recognition errors:
@@
To address these error types, we design rule-based strategies that use GLM-4 (GLM et al., 2024) to modify each positive caption and generate 10 hard negative examples. These text-based negatives are used in model training, and two representative types are shown in Figure 2.
3 MMGeoLM: Architecture and Training
@@
3.1 Architecture
We adopt the LLaVA architecture (Liu et al., 2024), which comprises three components: (1) a vision encoder, (2) a 2-layer MLP adapter, and (3) an LLM backbone. For the LLM backbone, we use MAmmoTH2-7B (Yue et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025). The vision encoder is based on AltCLIP (Chen et al., 2022b), configured with a maximum length of 512 tokens and a model size of 0.5B parameters.
3.2 Training Pipeline
@@
3.3 Hard Negative CLIP Training
In-batch Training. Conventional CLIP-trained vision encoders (Radford et al., 2021; Doveh et al., 2023b) typically combine random examples for a batch and adopt in-batch training loss over all the samples in each batch. For each batch containing samples $\{(I_{i},T_{i})_{i=1}^{N}\}$, where each $I_{i},T_{i}$ represents a pair of image and text caption, the loss is calculated as
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left[-\ln\frac{\exp(s(I_{i},T_{i}))}{\sum_{j}\exp(s(I_{i},T_{j}))}-\ln\frac{\exp(s(I_{i},T_{i}))}{\sum_{j}\exp(s(I_{j},T_{i}))}\right],$ (1)
where $\text{s}(\cdot,\cdot)$ denotes the similarity function over the feature encodings.
MMCLIP Training. In our data construction process, hard negative captions and images are generated for each image-caption pair without providing corresponding positive samples for these hard negatives. As a result, our MMCLIP training derives the loss solely from one positive sample and its associated hard negatives within each batch. This guides the model to discern subtle differences, enhancing its fine-grained geometric understanding. Formally, given a text-based negative training batch $\{I,T^{+},T_{i}^{-}|_{i=1}^{N}\}$, where $I$ represents the image, $T^{+}$ denotes the positive caption and $T_{i}^{-}$ is the negative caption conducted around the image. The loss function is formulated as follows:
$\mathcal{L}=-\ln\frac{\exp(s(I,T^{+}))}{\exp(s(I,T^{+}))+\sum_{i=1}^{N}\exp(s(I,T_{i}^{-}))}.$ (2)
Since negative samples are constructed around elements within the image, their score $s(I,T_{i}^{-})$ is expected to be higher than a randomly chosen one. From the above equation, we calculate the gradient for positive and negative samples:
$\frac{\partial\mathcal{L}}{\partial s^{+}}$ $=-\frac{\sum_{i=1}^{N}\exp(s^{-}_{i})}{\exp(s^{+})+\sum_{i=1}^{N}\exp(s^{-}_{i})}$
$\frac{\partial\mathcal{L}}{\partial s^{-}_{i}}$ $=\frac{\exp(s^{-}_{i})}{\exp(s^{+})+\sum_{i=1}^{N}\exp(s^{-}_{i})}$
where $s^{+}=s(I,T^{+})$ and $s_{i}^{-}=s(I,T_{i}^{-})$. Since the number of hard negatives is unrestricted, the gradient functions can be effectively optimized, leading to improved image-caption alignment. The training strategy is illustrated in Figure 2, where we also provide image-based negatives training batches $\{T,I^{+},I_{i}^{-}|_{i=1}^{M}\}$, the loss can be derived similarly to Equation 2. Unlike in-batch training, which relies on random batch-internal negatives, MMCLIP allows training with independently constructed negatives, making it suitable for unimodal negatives, such as rule-based or retrieval-based. We further validate different training strategies only for hard negatives in the ablation experiments.
4 Experiments
4.1 Main Experiment: MMGeoLM Evaluation
4.1.1 Datasets and Metrics
We assess the geometric reasoning capabilities of various LMMs across two types of benchmarks: multiple-choice and open-ended. The multiple-choice benchmark includes GeoQA (Chen et al., 2021), a geometric question answering task based on plane geometry; We-Math (Qiao et al., 2024), a visual mathematical reasoning task with questions of varying difficulty, solvable in two or three steps; and MathVista (Lu et al., 2023), widely used for evaluating LMMs’ performance. The open-ended benchmark is MM-Math (Sun et al., 2024), which features high discriminative difficulty sourced from secondary school-level problems. We use accuracy (ACC) as the metric. For We-Math, we assess multi-step problems (S2 and S3), while for MM-Math, we evaluate models across easy, medium, and hard categories using a test set of 700 problems. For MathVista, we evaluate geometry problem-solving (GEO) and algebraic (ALG) categories.
4.1.2 Baselines
The evaluated LMMs are categorized into two groups: Closed-source APIs: Claude-3-Opus (Anthropic, 2024), Claude-3.5-Sonnet, GPT-4o-20240513 (OpenAI, 2024a), and GPT-4V. Open-source LMMs: MAVIS-7B (Zhang et al., 2024c), G-LLaVA-7B (Gao et al., 2023), InternVL2-8B (Chen et al., 2024), LLaVA-OneVision-7B (Li et al., 2024), Qwen2.5-VL-7B-Instruct (Bai et al., 2025), Chimera-Reasoner-8B (Peng et al., 2024b), InternLM-XComposer2-7B (Dong et al., 2024), Phi-3-Vision-128K-Instruct (Abdin et al., 2024), and Math-LLaVA-13B (Shi et al., 2024). For the open-source category, we choose LMMs that have achieved strong results on geometric benchmarks (Chen et al., 2021; Lu et al., 2023). Additionally, we include human evaluation baselines, using scores from prior studies (Zhang et al., 2024c; Sun et al., 2024).
4.1.3 Overall Results
@@
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA (Peng et al., 2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-MAmmoTH2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-MAmmoTH2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-Math benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the w/o MMCLIP (In-batch trained AltCLIP with random negatives), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities. To rule out the possibility that performance gains stem from training–test leakage, we perform a contamination analysis (Appendix E) and we confirm that the evaluation sets contain no near-duplicate items.
4.2 Ablations
4.2.1 Ablation Study I: Different Negative Types